A simple hydraulic jack has a pump piston diameter 38.13 mm , and a load piston 190.65 mm . Calculate:


a) the effort force needed to lift a load of 1300 Newton


b) The distance moved by the load if the effort moves a distance 285

1
Expert's answer
2021-11-09T10:44:14-0500

"A_1 = \frac{\pi}{4}d^2 = \frac{\pi}{4}38.13^2=1141.88 mm^2"


"A_2 = \frac{\pi}{4}d^2 = \frac{\pi}{4}190.65^2=28547.19 mm^2"


a). The effort force needed to lift a load "F_1"


"\frac{F_1}{A_1}=\frac{F_2}{A_2}"


"F_1=\frac{F_2}{A_2}\times A_1"


"F_1=\frac{300}{28547.19}\times 1141.88 =11.99 N"


b) The distance moved by the load "l_2"


"F_1\times l_1=F_2\times l_2"


"l_2=\frac{F_1}{F_2}\times l_1"


"l_2=\frac{11.99}{1300}\times 285 =2.63 mm"